# A Fast Coordinate Descent Method for High-Dimensional Non-Negative Least Squares using a Unified Sparse Regression Framework

### v1

@@ 1 Introduction
Sparse regression methods have grown increasingly popular recently with the prevalance of high-dimensional data and large computing resource. Notably, the lasso [Tibshirani, 1996] is widely used for feature selection as it often sets many coefficients exactly to zero. It is particularly useful in ultra high-dimensional settings such as in genomics [He et al., 2022; Qian et al., 2020; Li et al., 2020] where feature selection is especially meaningful for interpretability. The non-negative least squares (NNLS) [Lawson and Hanson, 1995] is also a reliable tool for many practitioners across diverse fields as it often provides a physically meaningful and interpretable solution due to its empirical sparsity [Li and Speed, 2000; Slawski and Hein, 2013; Chou et al., 2022]. Some theoretical work exists around understanding this behavior [Bruckstein et al., 2008; Wang and Tang, 2009; Wang et al., 2011]. A natural extension of NNLS is the bounded variable least squares (BVLS) [Stark and Parker, 1995], which extends the non-negativity constraint to any box constraint. Similar sparsity occurs in high dimensions where many coefficients lie at their respective boundaries. The simplex constrained least squares (SLS) is at the heart of synthetic controls in causal inference [Abadie, 2021] where the potential outcome of no intervention for the treatment group is estimated using a weighted average of the donor pool’s outcomes. Abadie [2021] notes that we often observe sparsity in the weighted average, which is desirable for interpretability among practitioners especially when the donor pool size is large. Indeed, we see in Lei and Sudijono [2024][Table 2] that many real data examples have more donors than the pre-intervention time periods, which guarantees the existence of a sparse solution per our results in Section 3. Finally, the inequality simplex constrained least squares (ISLS) is commonly seen in portfolio optimization, e.g., a long-only portfolio under a Markowitz model with a leverage constraint so that the asset weights $w$ satisfy $w\geq 0$ and $\sum_{i}w_{i}\leq C$ for some $C>0$. The model may be in a high-dimensional setting if the covariance matrix of the assets $\Sigma\in\mathbb{R}^{p\times p}$ is further modeled by $k<p$ factors such that $\Sigma=FF^{\top}$ where $F\in\mathbb{R}^{p\times k}$. In this case, we often observe near zero weights for many assets using a standard interior point solver.
@@
While these methods may seem disparate at first glance, they share a deep connection through the more general problem of polyhedron constrained least squares. Our first contribution is a theoretical result in Section 3 that establishes the existence of a “sparse” solution to any polyhedron constrained least squares problem, thereby unifying our understanding of sparse solutions to the aforementioned methods. The notion of sparsity is extended to mean that the solution binds many linear inequality constraints as equality constraints in the polyhedron (see Definition 3.1). Moreover, we quantify the number of such binding constraints and show the local uniqueness property, which states that such sparse solutions are unique within the affine subspace defined by the binding constraints. We apply our results to NNLS, BVLS, SLS, ISLS, lasso, and box constrained lasso to show that the promised sparsity is at least $p-\operatorname*{rank}(X)+O(1)$ where $X\in\mathbb{R}^{n\times p}$ is the feature matrix. Lastly, we argue that the local uniqueness property helps explain the success of many active-set optimizers for these methods. Our second contribution is a computational result in Section 4 that uses the previous result as motivation to derive a coordinate descent based solver for BVLS and NNLS with screening and an active-set strategy. We show in Section 5 that our solver performs competitively in terms of accuracy and speed in high-dimensional settings.
@@ 3.2 Main results
We first restrict our attention to the simple case of a polyhedron $\mathcal{P}$ with a non-empty interior. Hence, the maximal affine constraints $S$ is empty. Suppose that $x\in\operatorname*{Hull}(Q,\mathcal{P})$ so that $x=Qw$ for some $w\in\mathcal{P}$. Our goal is to show that we can find a (possibly different) $\tilde{w}\in\mathcal{P}$ that binds at least a certain number of constraints of $A$ while reproducing $x$, i.e., $x=Q\tilde{w}$. Lemma 3.4 formalizes this claim and quantifies the number of binding constraints. Moreover, it also shows that the binding constraints form a full row rank sub-matrix of $A$, a fact that will become useful when we discuss the local uniqueness property in Theorem 3.8.
@@
Let $Q\in\mathbb{R}^{n\times p}$ be any matrix and $\mathcal{P}\subseteq\mathbb{R}^{p}$ be any polyhedron represented by $A\in\mathbb{R}^{m\times p}$ and $b\in\mathbb{R}^{m}$ with $S$ as its maximal affine constraints. Further let $V\in\mathbb{R}^{p\times\dim(\mathcal{P})}$ be any full column rank matrix such that $A_{S}V=\mathbf{0}$. Suppose that $\operatorname*{Null}(QV)\setminus\left\{\mathbf{0}\right\}\subseteq\operatorname*{Null}(A_{-S}V)^{c}$. If $x$ is in the $\mathcal{P}$-hull of $Q$, then $x$ can be written as a $\mathcal{P}$-combination of $Q$ with an element $w$ that binds at least $\min(\operatorname*{rank}(A)-\operatorname*{rank}(A_{S}),(\dim(\mathcal{P})-\operatorname*{rank}(QV))_{+})$ constraints of $A_{-S}$ that form a full row rank sub-matrix.
We give a proof of Theorem 3.5 in Section B.2. We first explain the assumptions of the theorem. As explained previously, a polyhedron may have an empty interior, and therefore, Lemma 3.4 cannot be applied directly. The remedy is the same as in the previous discussion; we first find the subspace in which the polyhedron lies (i.e., given by $S$) and remove the degrees of freedom from $Q$ and $\mathcal{P}$ to restrict to a lower-dimensional space. We then invoke Lemma 3.4 on the lower-dimensional space where the polyhedron is guaranteed to have a non-empty interior by Lemma A.5. The proof reveals that the new set of points is precisely $QV$ and the new polytope is represented by the constraint matrix $A_{-S}V$. Moreover, we show in the proof that $\operatorname*{rank}(A_{-S}V)=\operatorname*{rank}(A)-\operatorname*{rank}(A_{S})$. Lastly, we note that $\dim(\mathcal{P})-\operatorname*{rank}(QV)$ has been improved from $\dim(\mathcal{P})-n$. We argue that without loss of generality we may assume $QV$ to be full row rank since otherwise we can perform a QR-decomposition and use the “$R$” matrix in place of $QV$.
@@
Let $X\in\mathbb{R}^{n\times p}$ be a feature matrix and $y\in\mathbb{R}^{n}$ a response vector. Suppose $\mathcal{P}\subseteq\mathbb{R}^{p}$ is a polyhedron represented by $A\in\mathbb{R}^{m\times p}$ and $b\in\mathbb{R}^{m}$ that satisfies the conditions of Theorem 3.5 with $Q\equiv X$. Let $S$ be the maximal affine constraints of $\mathcal{P}$ and $V\in\mathbb{R}^{p\times\dim(\mathcal{P})}$ such that $A_{S}V=\mathbf{0}$. Consider the polyhedral regression given by 3. Then, there exists a solution $\hat{\beta}\in\mathbb{R}^{p}$ that binds at least $\min(\operatorname*{rank}(A)-\operatorname*{rank}(A_{S}),(\dim(\mathcal{P})-\operatorname*{rank}(XV))_{+})$ constraints of $A_{-S}$ that form a full row rank sub-matrix.
@@
Let $Q\in\mathbb{R}^{n\times p}$ be any matrix and $\mathcal{P}\subseteq\mathbb{R}^{p}$ be any polyhedron represented by $A\in\mathbb{R}^{m\times p}$ and $b\in\mathbb{R}^{m}$. Let $S$ be the maximal affine constraints of $\mathcal{P}$ and $V\in\mathbb{R}^{p\times\dim(\mathcal{P})}$ such that $A_{S}V=\mathbf{0}$. Let $T\subseteq[m-\left|S\right|]$ be any subset of size $(\dim(\mathcal{P})-\operatorname*{rank}(QV))_{+}$ such that $(A_{-S})_{T}$ forms a full row rank sub-matrix. Define $\mathcal{T}:=\left\{w\in\mathbb{R}^{p}:(A_{-S})_{T}w=(b_{-S})_{T}\right\}$. If $x$ is a $\mathcal{P}$-combination of $Q$ with an element $w\in\mathcal{P}\cap\mathcal{T}$, then $w$ is unique in $\mathcal{P}\cap\mathcal{T}$.
The local uniqueness property says that any solution $w$ in a sufficiently restricted subspace must be unique. We see then that once the conditions of Theorem 3.5 additionally hold, any such sparse solution must be locally unique given additionally the mild condition that
@@
or equivalently, $\operatorname*{rank}(A)-\operatorname*{rank}(A_{S})\geq\dim(\mathcal{P})-\operatorname*{rank}(QV)$. This condition formalizes the idea that there must be enough degrees of freedom in the set of points $QV$ and the constraint matrix $A$ combined for the existence of a solution in a sufficiently restricted subspace.
@@ 3.3.2 Bounded variable least squares
There exists a BVLS solution with at most $\operatorname*{rank}(X)$ non-binding values.
@@ Appendix B Proofs of the main results
We only detail the sketch of proof as the proof is nearly identical to the arguments made in Theorem 3.5.
As in Theorem 3.5, we may without loss of generality assume $Q$ is full row rank so that $n\equiv\operatorname*{rank}(Q)$. Moreover, we may further assume $\mathcal{P}$ has a non-empty interior since otherwise we may first restrict to the lower-dimensional space removing the maximal affine constraints by Lemma A.5. Hence, we assume $p\equiv\dim(\mathcal{P})$.
If $x$ is a $\mathcal{P}$-combination of $Q$ with an element $w\in\mathcal{P}\cap\mathcal{T}$, we may further restrict $w$ to a lower-dimensional subspace removing the affine constraints in $\mathcal{T}$ by Lemma A.5. Since $A_{T}$ has full row rank, we obtain $Q^{\prime}\in\mathbb{R}^{n\times k}$ and a solution $w^{\prime}\in\mathbb{R}^{k}$ such that $x=Q^{\prime}w^{\prime}$ where $k=p-(p-n)_{+}\leq n$. Since $Q^{\prime}$ must have full column rank, the solution for $w^{\prime}$ is unique, which establishes uniqueness for $w\in\mathcal{P}\cap\mathcal{T}$. ∎

### v2

@@ 1 Introduction
Sparse regression methods have grown increasingly popular recently with the prevalance of high-dimensional data and large computing resource. Notably, the lasso [Tibshirani, 1996] is widely used for feature selection as it often sets many coefficients exactly to zero. It is particularly useful in ultra high-dimensional settings such as in genomics [He et al., 2022; Qian et al., 2020; Li et al., 2020] where feature selection is especially meaningful for interpretability. The non-negative least squares (NNLS) [Lawson and Hanson, 1995] is also a reliable tool for many practitioners across diverse fields as it often provides a physically meaningful and interpretable solution due to its empirical sparsity [Li and Speed, 2000; Slawski and Hein, 2013; Chou et al., 2022]. Some theoretical work exists around understanding this behavior [Bruckstein et al., 2008; Wang and Tang, 2009; Wang et al., 2011]. A natural extension of NNLS is the bounded variable least squares (BVLS) [Stark and Parker, 1995], which extends the non-negativity constraint to any box constraint. Similar sparsity occurs in high dimensions where many coefficients lie at their respective boundaries. The simplex constrained least squares (SLS) is at the heart of synthetic controls in causal inference [Abadie, 2021] where the potential outcome of no intervention for the treatment group is estimated using a weighted average of the donor pool’s outcomes. Abadie [2021] notes that we often observe sparsity in the weighted average, which is desirable for interpretability among practitioners especially when the donor pool size is large. Indeed, we see in Lei and Sudijono [2024][Table 2] that many real data examples have more donors than the pre-intervention time periods, which guarantees the existence of a sparse solution per our results in Section 3. Finally, the inequality simplex constrained least squares (ISLS) is commonly seen in portfolio optimization, e.g., a long-only portfolio under a Markowitz model with a leverage constraint so that the asset weights $w$ satisfy $w\geq 0$ and $\sum_{i}w_{i}\leq C$ for some $C>0$. The model may be in a high-dimensional setting if the covariance matrix of the assets $\Sigma\in\mathbb{R}^{p\times p}$ admits a low-rank representation $\Sigma=FF^{\top}$ where $F\in\mathbb{R}^{p\times k}$ and $k<p$. In this case, we often observe near zero weights for many assets using a standard interior point solver.
@@
While these methods may seem disparate at first glance, they share a deep connection through the more general problem of polyhedron constrained least squares. Our first contribution is a theoretical result in Section 3 that establishes the existence of a “sparse” solution to any polyhedron constrained least squares problem, thereby unifying our understanding of sparse solutions to the aforementioned methods. The notion of sparsity is extended to mean that the solution binds many linear inequality constraints as equality constraints in the polyhedron (see Definition 3.1). Moreover, we quantify the number of such binding constraints and show the local uniqueness property, which states that such sparse solutions are unique within the affine subspace defined by the binding constraints. We apply our results to NNLS, BVLS, SLS, ISLS, lasso, and box constrained lasso to show that the promised sparsity is at least $p-\operatorname*{rank}(X)+O(1)$ where $X\in\mathbb{R}^{n\times p}$ is the feature matrix. Lastly, we argue that the local uniqueness property helps explain the success of many active-set optimizers for these problems. Our second contribution is a computational result in Section 4 that uses the previous result as motivation to derive a coordinate descent based solver for BVLS and NNLS with screening and an active-set strategy. We show in Section 5 that our solver performs competitively in terms of accuracy and speed in high-dimensional settings.
@@ 3.2 Main results
We first restrict our attention to the simple case of a polyhedron $\mathcal{P}$ with a non-empty interior. Hence, the maximal affine constraints $S$ is empty. Suppose that $x\in\operatorname*{Hull}(Q,\mathcal{P})$ so that $x=Qw$ for some $w\in\mathcal{P}$. Our goal is to show that we can find a (possibly different) $\tilde{w}\in\mathcal{P}$ that binds at least a certain number of constraints of $A$ while reproducing $x$, i.e., $x=Q\tilde{w}$. Lemma 3.4 formalizes this claim and quantifies the number of binding constraints. Moreover, it also shows that the binding constraints form a full row rank sub-matrix of $A$, a fact that will become useful when we prove the local uniqueness property in Theorem 3.8.
@@
Let $Q\in\mathbb{R}^{n\times p}$ be any matrix and $\mathcal{P}\subseteq\mathbb{R}^{p}$ be any polyhedron represented by $A\in\mathbb{R}^{m\times p}$ and $b\in\mathbb{R}^{m}$ with $S$ as its maximal affine constraints. Further let $V\in\mathbb{R}^{p\times\dim(\mathcal{P})}$ be any full column rank matrix such that $A_{S}V=\mathbf{0}$. Suppose that $\operatorname*{Null}(QV)\setminus\left\{\mathbf{0}\right\}\subseteq\operatorname*{Null}(A_{-S}V)^{c}$. If $x$ is in the $\mathcal{P}$-hull of $Q$, then $x$ can be written as a $\mathcal{P}$-combination of $Q$ with an element $w$ that binds at least $\min(\operatorname*{rank}(A)-\operatorname*{rank}(A_{S}),\dim(\mathcal{P})-\operatorname*{rank}(QV))$ constraints of $A_{-S}$ that form a full row rank sub-matrix.
We give a proof of Theorem 3.5 in Section B.2. We first explain the assumptions of the theorem. As explained previously, a polyhedron may have an empty interior, and therefore, Lemma 3.4 cannot be applied directly. The remedy is the same as in the previous discussion; we first find the subspace in which the polyhedron lies (i.e., given by $S$) and remove the degrees of freedom from $Q$ and $\mathcal{P}$ to restrict to a lower-dimensional space. We then invoke Lemma 3.4 on the lower-dimensional space where the polyhedron is guaranteed to have a non-empty interior by Lemma A.5. The proof reveals that the new set of points is precisely $QV$ and the new polyhedron is represented by the constraint matrix $A_{-S}V$. Moreover, we show in the proof that $\operatorname*{rank}(A_{-S}V)=\operatorname*{rank}(A)-\operatorname*{rank}(A_{S})$. Lastly, we note that $\dim(\mathcal{P})-\operatorname*{rank}(QV)$ has been improved from $\dim(\mathcal{P})-n$. We argue that without loss of generality we may assume $QV$ to be full row rank since otherwise we can perform a QR-decomposition and use the “$R$” matrix in place of $QV$.
@@
Let $X\in\mathbb{R}^{n\times p}$ be a feature matrix and $y\in\mathbb{R}^{n}$ a response vector. Suppose $\mathcal{P}\subseteq\mathbb{R}^{p}$ is a polyhedron represented by $A\in\mathbb{R}^{m\times p}$ and $b\in\mathbb{R}^{m}$ that satisfies the conditions of Theorem 3.5 with $Q\equiv X$. Let $S$ be the maximal affine constraints of $\mathcal{P}$ and $V\in\mathbb{R}^{p\times\dim(\mathcal{P})}$ such that $A_{S}V=\mathbf{0}$. Consider the polyhedral regression given by 3. Then, there exists a solution $\hat{\beta}\in\mathbb{R}^{p}$ that binds at least $\min(\operatorname*{rank}(A)-\operatorname*{rank}(A_{S}),\dim(\mathcal{P})-\operatorname*{rank}(XV))$ constraints of $A_{-S}$ that form a full row rank sub-matrix.
@@
Consider the setting of Theorem 3.5. Suppose that
$\operatorname*{rank}(A)-\operatorname*{rank}(A_{S})\geq\dim(\mathcal{P})-\operatorname*{rank}(QV)$
Let $w$ be the sparse representation that binds at least $\dim(\mathcal{P})-\operatorname*{rank}(QV)$ constraints of $A_{-S}$. Let $T$ be the set of indices such that $(A_{-S})_{T}$ denotes the binding constraints. Define $\mathcal{T}:=\left\{w\in\mathbb{R}^{p}:(A_{-S})_{T}w=(b_{-S})_{T}\right\}$. Then, $w$ is the unique element in $\mathcal{P}\cap\mathcal{T}$ such that $x=Qw$.
The local uniqueness property says that the sparse solution $w$ given by Theorem 3.5 must be unique in a sufficiently restricted subspace. The condition for Theorem 3.8 can be equivalently written as
@@
This mild condition formalizes the idea that there must be enough degrees of freedom in the set of points $QV$ and the constraint matrix $A$ combined for the existence of a solution in such a sufficiently restricted subspace. In many practical settings, this condition will hold true as we often have $\operatorname*{rank}(A)\geq p$.
@@ 3.3.2 Bounded variable least squares
There exists a BVLS solution with at most $\operatorname*{rank}(X)$ non-boundary values.
@@ Appendix B Proofs of the main results
We may assume without loss of generality $\mathcal{P}$ has a non-empty interior since otherwise we may first restrict to the lower-dimensional space removing the maximal affine constraints by Lemma A.5. Hence, we assume $p\equiv\dim(\mathcal{P})$. We may further assume $Q$ is full row rank so that $n\equiv\operatorname*{rank}(Q)$. We are now in the setting of Lemma 3.4 where $Q$ is full row rank and $\mathcal{P}$ has a non-empty interior.
Consider the proof of Lemma 3.4. Recall that the proof procedure is to repeatedly find directions $v\in\operatorname*{Null}(Q)$ such that $w=y+tv$ binds a constraint (where $x=Qy$), say $a_{1}$. Let $V\in\mathbb{R}^{p\times(p-1)}$ be as in Lemma A.5 that reduces the polyhedron by one dimension, namely, restricting to the affine space $\left\{w\in\mathbb{R}^{p}:a_{1}^{\top}w=b_{1}\right\}$. We claim that $\operatorname*{rank}(QV)=\operatorname*{rank}(Q)$. Indeed, it is equivalent to show that $\dim(\operatorname*{Null}(QV))=\dim(\operatorname*{Null}(Q))-1$. This is easily seen by the fact that since $a_{1}^{\top}v\neq 0$ by Lemma A.4 and $V^{\top}a_{1}=\mathbf{0}$, we must have $v\notin\operatorname*{span}(V)$. Since the rank is preserved as we apply induction, we see that eventually $Q$ will be reduced to an invertible matrix, in which case, the representation for $x$ is unique. ∎